A kilometer is about 5/8 mile. About how many miles are in 4 2/5? What is the proportion and what is the correct answer?
Answer by sachi(548)

A kilometer is about 5/8 mile.
4 2/5=(4*5+2)/5=22/5
no of miles=(22/5)*(5/8)=11/4